# On the Importance of Embedding Norms in Self-Supervised Learning

### v1

@@ C.1 Experimental Setup
Due to computational constraints, we run with batch-size 256 in SimCLR. Although each batch is still 256 samples in SimSiam, we simulate larger batch sizes using gradient accumulation. Thus, our default batch-size for SimSiam is 1024.
@@ C.2 Opposite-Halves Effects
Table S2: $k$-nn accuracies for SimSiam trained with various batch sizes. We performed training for both the default and cut-initialized variants and reported $k$-nn accuracies at 100 and 500 epochs.
In accordance with this, Table S1 shows that, after one epoch of training, embeddings have angle greater than $\pi/2$ at a rate of around $5\%$ and $25\%$ for SimCLR and SimSiam/BYOL, respectively. So even if the ‘strongest’ variant of the opposite-halves effect is not occurring, a weaker one may still be. However, very early into training (epoch 16), every method has a rate of effectively 0 for the opposite-halves effect. Furthermore, the rates in Table LABEL:tbl:opposite_half_effect measure how often $\phi_{ij}>\frac{\pi}{2}$. This is the absolute weakest version of the opposite-halves effect. Thus, while some weak variant of the opposite-halves effect may occur at the beginning of training, it does not have a strong impact on the convergence dynamics and, in either case, disappears quite quickly.
@@ C.3 Weight Decay
We evaluate the effect of weight decay in the imbalanced setting in S3, which is an analog of Figure 5 for the imbalanced Cifar-10 dataset detailed in Section 6. We again see that using weight decay controls for the embedding norms and improves the convergence of both models. In correspondence with the other results on imbalanced training, we find that stronger control over the embedding norms leads to improved convergence: the high weight decay value does not perform as poorly on SimCLR as in Figure 5 and, on SimSiam, outperforms the other weight decay options.
@@
Figure S3: An analog to Figure 5 performed on the exponentially imbalanced Cifar-10 dataset. Weight decays are [00, $1$e-$5$, $5$e-$2$] for SimCLR and [00, $5$e-$4$, $5$e-$2$] for SimSiam. We plot the effective learning rate in the bottom row, calculated in accordance with Section 6.
@@ C.4 Cut-Initialization
Figure S4: The effect of cut-initialization on Cifar10 SSL representations. $x$-axis and embedding norm’s $y$-axis are log-scale. $\lambda=5$e44-4- 4 in all experiments.
@@ Appendix D More details on gradient scaling layer
An implementation of our GradScale layer can be found in Listing 1. We note that this layer is purely a PyTorch optimization trick and does not amount to implicitly choosing a different loss function:

### v2

@@ C.1 Experimental Setup
Due to computational constraints, we run with batch-size 256 in SimCLR. Although each batch is still 256 samples in SimSiam, we simulate larger batch sizes using gradient accumulation. Thus, our default batch-size for SimSiam is 1024. Our base learning rate is set to $0.18\cdot\text{BatchSize}/256$ for SimCLR and $0.12\cdot\text{BatchSize}/256$ for SimSiam. We employ a 10-epoch linear warmup followed by cosine scaling. For all cases, we use $k=200$ for the $k$-NN classifier and apply it on the normalized embeddings.
@@ C.2 Opposite-Halves Effects
Table S2: $k$NN accuracies for SimSiam trained with various batch sizes. We performed training for both the default and cut-initialized variants and reported $k$NN accuracies at 100 and 500 epochs.
@@
In accordance with this, Table S1 shows that, after one epoch of training, positive pairs of embeddings have angle greater than $\pi/2$ at a rate of around $5\%$ and $25\%$ for SimCLR and SimSiam/BYOL, respectively. So even if the ‘strongest’ variant of the opposite-halves effect is not occurring, a weaker one may still be. However, very early into training (epoch 16), every method has a rate of effectively 0 for the opposite-halves effect. Furthermore, the rates in Table S1 measure how often $\phi_{ij}>\frac{\pi}{2}$. This is the absolute weakest version of the opposite-halves effect. Thus, while some weak variant of the opposite-halves effect may occur at the beginning of training, it does not have a strong impact on the convergence dynamics and, in either case, disappears quite quickly.
@@ C.3 Weight Decay
Table S3: Linear probe accuracies at epoch 500 for default, cut-initialized and GradScale training on a subset of our image datasets.
Cifar-100 Tiny Imagenet
SimCLR Default 59.8 41.9
Cut ($c=3$) 63.2 42.8
GradScale 62.2 43.2
SimSiam Default 63.7
Cut ($c=9$) 64.2
Table S4: Final embedding norms at epoch 100 for SimCLR trained with Adam optimizer on Cifar-10, demonstrating the effect of different mitigation strategies on embedding magnitude.
Method Embedding Norm
Default 81.0
Cut 2.1
GradScale 174.8
We evaluate the effect of weight decay in the imbalanced setting in Figure S3, which is an analog of Figure 5 for the imbalanced Cifar-10 dataset detailed in Section 6. We again see that using weight decay controls for the embedding norms and improves the convergence of both models. In correspondence with the other results on imbalanced training, we find that stronger control over the embedding norms leads to improved convergence: the high weight decay value does not perform as poorly on SimCLR as in Figure 5 and, on SimSiam, outperforms the other weight decay options.
@@
Figure S3: An analog to Figure 5 performed on the exponentially imbalanced Cifar-10 dataset. Weight decays are [00, $1$e-$5$, $5$e-$2$] for SimCLR and [00, $5$e-$4$, $5$e-$2$] for SimSiam. We plot the effective learning rate in the bottom row. This is calculated by scaling the learning rate by the inverse of the mean embedding norm.
@@ C.4 Cut-Initialization
Figure S4: The effect of cut-initialization on Cifar-10 SSL representations. $x$-axis and embedding norm’s $y$-axis are log-scale. $\lambda=5$e44-4- 4 in all experiments.
@@ Appendix D More details on gradient scaling layer
An implementation of our GradScale layer can be found in Listing LABEL:alg:grad_scaling and pseudocode in Algorithm 1. We note that this layer is purely a PyTorch optimization trick and does not amount to implicitly choosing a different loss function: